[Hallc_running] [New Logentry] Saturday RC update for Hall C

mack at jlab.org mack at jlab.org
Sat Jun 29 12:00:07 EDT 2019

Logentry Text:
We have some shift leader gaps on owl during the coming week, although we have TO's and 3rds. Field commissions of 3rd-->SL might not always be appropriate, so please consider signing up.

The last day was a mix of beam off due to accelerator problems and a Hall C angle change requiring spotters. Otherwise, production.

It's probably worth having an RC meeting at 4:30pm so we can see where the run plan stands wrt the change to 4-pass.

TMI follows:

FRI DAY shift - The Q2 = 0.425,W=2.20 , theta_pi = 5.75deg setting was essentially completed.** Stephen posted plots of accumulated statistics at https://logbooks.jlab.org/entry/3696076 . Regarding potentially low (e,e'pi) yields, there was some discussion of cuts changes by Dave G:  https://logbooks.jlab.org/entry/3696035  , and Simona had left a great A/C party to work all nite and demonstrate that the ELREAL trigger thresholds were fine https://logbooks.jlab.org/entry/3695871 . See Tanja's day shift summary at https://logbooks.jlab.org/entry/3696053 . **Despite the perfection of the run plan, it's easy to miss the block of Al dummy running at the bottom of the page, particularly during the chaos of staging an angle change. Crews are now doing the Al dummy first thing after an angle change.

FRI SWING shift - The SHMS angle was changed to 7.75deg by a patient team of spotters. Deepak posted his very satisfactory results from an Unser wire calibration taken during a recent beam off period https://logbooks.jlab.org/entry/3696051 . After hours of no beam due to the inability to properly cycle the 3C arc magnets, and no end in sight, we (MCC/PS tech/Hall C) agreed  it would be expedient to improperly cycle the magnets at lower Imax if this would restore beam. The downside is that we'll have to determine the 3-pass beam energy by indirect methods; hopefully the PS repair can take place opportunistically when we get to 4-pass. See Garth's swing shift summary at https://logbooks.jlab.org/entry/3696063 .

SAT OWL shift - The crew completed the ongoing SHMS angle setting, then changed it remotely to 9.74deg and continued to take data. I don't want to jinx us, but the crew had an ood-gay ift-shay. See Mireille's owl shift summary at https://logbooks.jlab.org/entry/3696289 .

SAT DAY shift - The crew completed the 9.74deg settings, and Markus just reported they are now in a flurry of angle and momentum settings to continue (e,e'pi) production on "page 10". As I write, looks like I did jinx us. Asking for Controlled Access .... 

Still on the to-do list for 3-pass:
- H elastic singles: p=-2,3, -1,96, -1.85
- heep check coin run
DROPPED:  Arc energy measurement  



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3696488
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190629/1495e94f/attachment.html>

More information about the Hallc_running mailing list